OpenStudy (anonymous):

Becca family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last?

OpenStudy (anonymous):

What you have to do is 4/(5/8) This can be solved by taking the inverse of 5/8 8(4)/5 = 32/5 equals a little more than 6 days

OpenStudy (jdoe0001):

\(\bf \cfrac{4}{\frac{5}{8}}\implies \cfrac{\frac{4}{1}}{\frac{5}{8}}\\ \quad \\ \textit{recall that }\cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\qquad thus\\ \quad \\ \cfrac{\frac{4}{1}}{\frac{5}{8}}\implies \cfrac{4}{1}\cdot \cfrac{8}{5}\)